the sum of two angles is 116 and their difference is 24 .find the measure of each angle of the triangle

Answer by kaushik31(2)

let two anles be x and y
by the problem,
x+y=116............(1)
x-y=24..............(2)
adding the equations we get,
x+y+x-y=116+24=140
2x=140
x=70.........(3)
putting the value of x in equation (2) we get,
x-y=24
70-y=24
y=70-24
y=46
ans:- x=70 y=46